# Parallelism Meets Adaptiveness: Scalable Documents Understanding in Multi-Agent LLM Systems

### v2

@@ 2 Related Work
Multi-agent systems built on large language models (LLMs) have rapidly emerged as a powerful paradigm for solving complex, multi-stage tasks. Early frameworks such as AutoGPT and BabyAGI introduced single-agent recursive task planners with memory and subtask management, though they were limited by brittle, linear execution flows and lacked true inter-agent coordination [torres2023auto, babyagi2023github]. Subsequent systems like CAMEL [zhu2023camel] and CrewAI [crewai2024github] introduced role-based delegation and dialogue-based collaboration among agents. LangGraph formalized agent workflows using graph structures [langgraph2024], but did not incorporate mechanisms for feedback loops or task reassignment.
Recent meta-agent and hierarchical systems, such as MetaGPT [hong2023metagpt] and Voyager [xu2023voyager], demonstrate improved planning, code synthesis, and environmental interaction. However, they typically assume static roles or unidirectional flows of control, lacking adaptiveness in task routing or feedback integration. The HuggingGPT framework [shen2023hugginggpt] explores task orchestration by using ChatGPT to coordinate models on Hugging Face, yet still relies on fixed assignments. Generative Agents [park2023generative] illustrate long-term memory and simulation of social behaviors, providing inspiration for agent autonomy but not competitive refinement.
A more recent trend incorporates reflective or competitive coordination. For instance, systems like ChatDev [qian2024chatdev] and AgentVerse [chen2024agentverse] explore inter-agent negotiation and emergent behaviors. Others such as GameGPT [chen2023gamegpt] and DesignGPT [ding2023designgpt] use LLM agents in creative or design-oriented domains, but often lack deep feedback-driven revision loops. Reflective multi-agent systems [bo2024reflective] combine agent memory with iterative critique, approaching the kind of meta-level coordination our system emphasizes.
Our work is also informed by broader surveys of multi-agent LLM collaboration [wang2023survey, lamalfa2025miss], which identify key challenges including error propagation, brittle delegation, and a lack of robustness under dynamic task flows. In contrast, our framework explicitly introduces dynamic routing, feedback-based revision, shared long-term memory, role self-optimization, and competitive evaluation—yielding a system that is not only collaborative but also adaptively reflective.
In a complementary domain, Wu et al. [wu2025warehouse] explore competitive task scheduling for warehouse robots using reinforcement learning, demonstrating that concurrent multi-agent evaluation can improve efficiency even in physical task settings. Cross-domain applications of multi-perspective analysis and feedback-driven reconstruction, such as in civil infrastructure using deep learning and multiview stereo methods [dan2024evaluation], further support the value of evaluation-oriented designs in complex systems. We build on this insight in the LLM space by introducing competitive agent evaluation to select the best output from parallel attempts.
@@ 5.3 Modularity and Extensibility
Figure 1: System architecture for adaptive coordination and competitive evaluation in multi-agent LLM systems. Solid lines represent data flow and task assignment; dashed lines represent feedback loops and evaluator interventions.
@@ 6.1 Example Prompt and Comparative Outputs
Figure 2: Comparative system outputs for a compliance query on off-balance sheet arrangements. The static system failed to detect the disclosure; the adaptive system surfaced a partial statement; the full system accurately identified and contextualized the \$150 million receivables securitization, aligning with the gold-standard reference.

### v3

@@ 2 Related Work
Multi-agent systems built on large language models (LLMs) have rapidly emerged as a powerful paradigm for solving complex, multi-stage tasks. Early frameworks such as AutoGPT and BabyAGI introduced single-agent recursive task planners with memory and subtask management, though they were limited by brittle, linear execution flows and lacked true inter-agent coordination [13, 17]. Subsequent systems like CAMEL [18] and CrewAI [11] introduced role-based delegation and dialogue-based collaboration among agents. LangGraph formalized agent workflows using graph structures [12], but did not incorporate mechanisms for feedback loops or task reassignment.
Recent meta-agent and hierarchical systems, such as MetaGPT [6] and Voyager [16], demonstrate improved planning, code synthesis, and environmental interaction. However, they typically assume static roles or unidirectional flows of control, lacking adaptiveness in task routing or feedback integration. The HuggingGPT framework [10] explores task orchestration by using ChatGPT to coordinate models on Hugging Face, yet still relies on fixed assignments. Generative Agents [8] illustrate long-term memory and simulation of social behaviors, providing inspiration for agent autonomy but not competitive refinement.
A more recent trend incorporates reflective or competitive coordination. For instance, systems like ChatDev [9] and AgentVerse [3] explore inter-agent negotiation and emergent behaviors. Others such as GameGPT [2] and DesignGPT [5] use LLM agents in creative or design-oriented domains, but often lack deep feedback-driven revision loops. Reflective multi-agent systems [1] combine agent memory with iterative critique, approaching the kind of meta-level coordination our system emphasizes.
Our work is also informed by broader surveys of multi-agent LLM collaboration [14, 7], which identify key challenges including error propagation, brittle delegation, and a lack of robustness under dynamic task flows. In contrast, our framework explicitly introduces dynamic routing, feedback-based revision, shared long-term memory, role self-optimization, and competitive evaluation—yielding a system that is not only collaborative but also adaptively reflective.
In a complementary domain, Wu et al. [15] explore competitive task scheduling for warehouse robots using reinforcement learning, demonstrating that concurrent multi-agent evaluation can improve efficiency even in physical task settings. Cross-domain applications of multi-perspective analysis and feedback-driven reconstruction, such as in civil infrastructure using deep learning and multiview stereo methods [4], further support the value of evaluation-oriented designs in complex systems. We build on this insight in the LLM space by introducing competitive agent evaluation to select the best output from parallel attempts.
@@ 5.3 Modularity and Extensibility
Figure 1: System architecture for adaptive coordination and competitive evaluation. The framework integrates dynamic task routing with a parallel execution module for high-ambiguity tasks. Solid lines represent data flow and task assignment; dashed lines represent feedback loops and evaluator interventions.
@@ 6.1 Example Prompt and Comparative Outputs
Figure 2: Qualitative comparison on a high-ambiguity compliance task. We query the systems regarding “off-balance sheet arrangements.” While the Static baseline misses the disclosure and the Adaptive system offers only a partial summary, the Full Framework correctly identifies and quantifies the \$150 million impact, matching the gold-standard reference.